A population consists of the values (1, 3, 4). If the samples of size 2 will be drawn from the population, find the variance of the sampling distribution of the means.

1
Expert's answer
2022-04-21T08:31:49-0400

The population mean:

"\mu=\cfrac{1+3+4}{3}=\cfrac{8}{3}."


The population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 1-\cfrac{8}{3},3-\cfrac{8}{3},4-\cfrac{8}{3}
\end{Bmatrix}="

"=\begin{Bmatrix}
-\cfrac{5}{3}, \cfrac{1}{3},\cfrac{4}{3}
\end{Bmatrix},"

"\sigma^2=\bigg(\cfrac{-5}{3}\bigg)^2\cdot \cfrac{1}{3}+\bigg(\cfrac{1}{3}\bigg)^2\cdot \cfrac{1}{3}+\bigg(\cfrac{4}{3}\bigg)^2\cdot \cfrac{1}{3}=\\
=\cfrac{42}{27}=\cfrac{14}{9}=1.556."


The variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{1.556}{2}=0.778."